OpenStudy (anonymous):

Solve the following equation x/2 +x/3 =5

OpenStudy (anonymous):

factor x

OpenStudy (anonymous):

x/2+x/3=5 3*x/3*2+x*2/3*2=5 (3x+2x)/6=5 5x=5*6 5x=30 x=6

OpenStudy (anonymous):

@arch1 please don't do the work for them!

OpenStudy (anonymous):

@wired okay